Meeting Information

  • 1930 UT (1330 MDT)
  • AOC 317; IP:; phone: +1 575-835-7370



  • 20 Hz backed out; still undergoing testing
  • Brief foray into wide band modes (Michael)
  • Multi-config mostly working; not yet ready for prime time (Michael)
  • Staff Travel:
    • Michael: 27 Apr - 4 May (i.e., leaves tomorrow.)
    • Lorant: 9 May - 22 May
    • Joan: 13 May - 18 May
    • Joe: 19 - 24 May
    • Gustaaf: 1 - 4 June

Commissioning review/reports

Notes for week of 05 Apr 2010 (Black: old but still relevant; Blue: new)

  • Meta summary

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
    OST not ready (done manually) OPT selected limits not possible - output averaged SBs affected by zero visibilities - needs check on short pol'n runs - can't handle mixed configs (no X band with other than Ka < 32 GHz)  

  • Went to a new system for the status of the OSRO observing/priorities. See:
  • Priorities notes (Commissioning, OSRO, RSRO, Demo science) (Note: Not changed from last week due to some set backs in the system.).
    • Key testing:
      • Multi-configuration SBs
      • Push to wider bandwidths
      • Switched power implementation

  • Quick Software Systems review:
    • OPT (Lorant): Waiting to get the production version updated. Switch after testing on Tuesday. L band frequencies updated. For staff, use production.
    • Executor (Barry): -
    • Scheduling (Joan, Bryan): Planning for more OST tests once the multi-config testing has been completed.
    • CASA (Juergen): 3.01.1 released last Thursday. 3.0.2 updated (better pol'n angle correction); linux versions are automatically updated; Mac users must do manually. Commissioners should be using the casapy-test version. Gustaaf noted that he had requested that the latest version of importevla not be put into the casapy-test as it hadn't been tested and is used by the archive system. Currently this is broken as the exportuvfits is failing. Discussions to handle this will follow this meeting.
    • AIPS (Eric): REWAY saga: hunted down much of the issue in bad flagging/bad data. Eric will report back on other issues identified by Frazer. Frazer noted that the weights on each antenna can wander by 50%. Mislabeled times in WEATHER table; need to be defensive here (e.g., provide a warning if absurd opacity is generated).
  • Hardware notes:
    • EA12 passed operations testing; ptg/collimation tests to be completed; moves on 22 Apr; Rx outfitting for L, Ku,K and Q by end of next week tentatively.
    • EA06 will be removed from the array for further work.
    • Notes from 4/23 meeting with engineers:
      • Joe: Many issues recently not readily tied to hardware (e.g., missing scans, missing BDFs, zeroes issue, etc).
      • Joe: Still having issues in the reporting; generally people are noting what they flagged but not why; we need to provide further guidance on this at Monday's meeting: ACTION: JOE - provide guidelines for reporting.
      • Peggy: Noted that the maintenance day coordination across all groups could be improved, particularly with regard to software changes (which often aren't expected to disrupt the system but sometimes do). ACTION: PEGGY - work with Bryan to better understand the full activity on the system each maintenance day.
      • Peggy: Would like to see the stress tests reinstated (see above). Reiterated by the engineers who were used to have this first line of defense when returning the system in good conscience.
      • Joe: Work through the most egregious antennas:
        • EA06: Issues with amplitude changes confirmed (seen at input/output to T304. Worked on Wednesday - would like confirmation that the issue still exists: ACTION: JOE - generate ticket to review data after 22 Apr looking for the 57.6 second period amplitude variation as well as bandpass edge variations.
        • EA05: Issues with the jumps at edges has been handled by inserting a 12" cable to the synthesizer (in front of splitter). This will be done tonight to EA05 and Joe will be notified. ACTION: JIM - have cable installed in EA05 ACTION: JOE - review data Saturday to look for resolution.
        • EA03: synthesizer was changed out; need to confirm that it has solved the problem. ACTION: JOE - confirm/deny phase jumps at scan boundaries.
        • EA20: Amplitude jumps could also be a servo issue; to be reviewed; Jim also suggested this could be the FRM.
        • EA23: Recent issue with amplitude drops consistent with a pointing drift. Recent work done on servo thought had resolved this (fixed slow source slew). ACTION: JIM - will investigate this.
        • EA26: Jim wanted to confirm that there were no jumps (B/D only). ACTION: JOE to confirm and send note to Jim.
        • EA28: Jim has a note about no fringes for this antenna. ACTION: JOE will confirm and send note to Jim.

OSRO Priority Testing Status

System Verification

  • Array/Receivers (Rick)
    • EA12 has joined the array. (get array location from Gene)
  • WIDAR update (Michael)
    • Multi-config/'hands-off' observing status
    • Narrow band status: Looks good (in use now). Still hunting down the zeros issues.
    • CRM checks are ongoing; some failures in this to be hunted down.
    • Wideband checkout status
    • Restricted testing this week
    • Issues:
      • Missing scans issues; being reviewed.
      • Lost fringes in the middle of scans (EA03 - went away and came back; EA01 - 5 seconds of fringe loss); review ECDATA-160

Band Checkouts (ECDATA-55)

  • P Band Status: %Red%Testing
    • Antennas with no P-band Rx: 7, 16, 22, 24 (no feed); 1, 4, 5 Rxs are not working.
  • L Band Status: Ready
    • Frequency labeling (ECDATA-9, ECDATA-10): Ok
    • Sky orientation (ECDATA-10): Ok
    • Gain stability (ECDATA-23): Ok; some bandpass phase evolution seen; follow-up on really blank sky.
    • Tuning: updated.
    • Antennas with no L-band: EA06, EA20, EA26
    • Integrating down (ECDATA-23):
      • Averaging in time is ok; averaging between subbands has problems.
      • L band (really) blank field (ECDATA-40)
    • L band, HI checkout (ECDATA-47)

  • S Band Status: Ready.
    • S band edge tests (similar result to C for tunings near 4 GHz; potentially some satellite downlink RFI; review of aliasing/out-of-band intermodulation is under review (Rick)
    • Frequency labeling (ECDATA-24): Ok
    • Sky orientation (ECDATA-18): Ok
    • Gain stability (ECDATA-24): Ok (no jumps seen)
    • Tuning (ECDATA-24): Reviewing default frequency based on birdies; recommendation to shift to multiple of 128 MHz .
    • Integrating down (ECDATA-18): Ok
    • Antennas with S band: EA08, EA09, EA15, EA22, EA24, EA27, EA28

  • C Status: Ready
    • Frequency labeling: Ok. Will get more information from Eric's analysis.
    • Sky orientation (ECDATA-35): Ok
    • Gain stability (ECDATA-35): Ok
    • Tuning: Low frequency tuning issue; stepped up for last AH1004.
    • Integrating down (ECDATA-25):
    • Antennas with no C-band: EA13

  • X Band Status: Ready
    • Frequency labeling : ECDATA-17): Ok
    • Sky orientation (ECDATA-17): Ok
    • Gain stability (ECDATA-26): Ok (no jumps)
    • Tuning: No issues
    • Integrating down (ECDATA-26): Ok
    • Antennas
    • RFI: none seen

  • K Band Status: Ready

  • Q Band Status: Ready.
    • Need Q band suite of tests.
    • Antennas with no Q-band: EA10, EA12, EA17

Mode testing

  • Polarimetry Status:
    • C (ECDATA-12): Ready
    • L (ECDATA-42): Ready
    • Need High Frequency Test
    • Eric reviewing AH1004 data for AIPS
    • Rick: Pol'n tests (receiver rotation) to be scheduled

  • Wideband L (AB1345 - ) Sanjay
  • Solar (Rick) status and path forward. Working with L, Ku (special Rx setups in these two bands). TP tests coming.
  • Phased Array: Adam+ did observations of a maser line; mostly worked.

OSRO Proposal Testing (ECDATA-34)

Commissioning plans for the week

System Tests

  • [Michael]: Enable multi-configuration SBs (with no hand-editing)
  • [Joan/Michael/Bryan]: Enable Dynamic Scheduling (diagnostics/debugging)
  • [Rick/Michael]: RSRO Observing Modes
  • Area Commissioning Tests
    • Group Leads should post plans to their respective JIRA tickets.
    • When SBs are developed, notify: Joe, Joan, Michael for prioritization in week's scheduling
    • Current priorities are:
      • WIDAR commissioning (driven by Michael); [TCOR; TOSR0001]
      • End-to-end or "Hand's off" observing (driven by Bryan)
      • General System Verification (driven by Rick)
      • Area Commissioning
        • 1. TOSR
        • 2. TSPE
        • 3. THIG/TMID/TLOW
        • 4. TPOL
        • 5. TSCH
        • ---
        • TCAL: plan proposed; needs review and scheduling.
        • TPLA/TSOL will need vetting soon. Need planned set of observations.
        • TDEM is complete for summer school (may still want mosaic for continuum polarimetry); other demo science will be done as filler projects.
        • TRAP need plan; currently ToO being handled via scheduling officer.

Area Test D config time required (hrs) SBs submitted LST Start range
SV * 10 -  
SV: Correlator * 20 TOSR0001*  
Observing Mode: Polarization High Frequency test 7 TPOL0001
Observing Mode: Spectral Line radio recomb line tests
3C84 bandpass tests
TSPE0006 (sb1311164)
TSPE0005 (sb1261338)
Observing Mode: Summer School IRC+10216 (spectral line)
  TDEM0003 (sb1345754)
TDEM0001 (sb1218006)
Observing Mode: Demo Science IRC+10216 spectral sweep      
Observing Support: Calibrator survey/models Two frequencies within each band (L/C/K/Ka/Q) for 4 sources (3C48, 3C147, 3C286, 3C138) 20 Needed; needs multi-config Not-yet-scheduled
Observing Support: Robustness Stress test observation (multi-band) 30 min/per diagnostic TOSR0005 Not-yet-scheduled
Algorithm Development wideband, wide field imaging
A-projection test
RSRO test


TALG0002 (sb1179032)
TALG0003 (sb1351545)

  • *Need any further D config time requests by 5/3*

OSRO project tests

  • Several projects released; metric is 70% observing efficiency (55% for fast switched).
  • Tool from Lorant to see available SBs: /home/mchost/evla/bin/sbStatReport
  • Priorities at:
  • List of known problems (can be used by OSRO commissioning team (van Moorsel)
      • Med Priority [Rupen]: X, Ku band subbands are inverted (IF 1 is labelled as IF4, IF2 as 3, IF3 as 2, IF 4 as 1). Noted that this had been tested but there were enough changes to warrant a re-do.
      • Med Priority [Butler]: Scan labeling synchronization. Any new instances seen? None reported
      • Med Priority [Perley]: Issues integrating down when averaging by channel. Okay on blank fields.
      • Med Priority [TBD]: Subband 0 problems: Being reviewed on Wednesday morning at the WIDAR meeting. Butler to report.
      • Med Priority [TBD]: Phase offsets within scans (phase jump at scan boundaries). R&L phase offset on some scans - probably L302 LO problem, needs tracking down. MR: Seen on EA07, EA25 (both RR and LL)
      • Med Priority [TBD]: Added. Phase change when frequency changes (needs to be tested).
      • Med Priority [TBD]: Anti-aliasing does not behave as advertised.
      • Low Priority [TBD]: Occasional zero-amplitude visibilities (needs software fix - what level). UVLOD and FITLD can handle this in AIPS; flagdata clip is effective in CASA; still need to understand and fix at the source.
      • Low Priority [TBD]: Occasional pure-noise scans.
      • Low Priority [Rupen/Pokorny]: Delay clunking; software fix pending.
      • Low Priority [TBD]: First scan often bad.
      • Low Priority [TBD]: Phase (and amp) wobbles; long history.


  • Observations
  • Residents
    • Manuel Aravena (POC: Chris Carilli); area(s): High frequency checkout (reference pointing tests)
    • Crystal Brogan (POC: Claire Chandler); area(s): High frequency calibration
    • Laura Chomiuk (POC: Miller Goss; area(s): Correlator checkout; Spectral line observing)
    • Todd Hunter (POC: Claire Chandler; area(S): High frequency checkout
    • James Miller-Jones (POC: Claire Chandler; area(s): Demo Science)
    • Jacqueline van Gorkom (POC: Miller Goss; area(s): Correlator checkout; Spectral line observing)

Demo Science : ECDATA-153

  • No further D config observations


  • [Rupen]: Develop daily/post-maintenance day checklists for the system. To be completed as part of stress tests.
  • [McMullin]: Post Eric's AIPS Guide as template for reduction strategy. Working on how to format this; twiki not good for providing a polished look to users. Want to use mediawiki; discussions with CV ongoing, to be resolved today.
  • [McMullin]: Add ECSO Projects to JIRA/tracking.
  • [McMullin]: Review/recover metrics for Priority 0 items.
  • [Perley/Dhawan]: Develop test to develop/demonstrate readiness for high frequency observing.
  • [Perley]: Develop set of frequency accuracy tests (X band is complete).
  • [Perley]: Perform a frequency sweep test at different bands to provide to engineers for review (Q2).
  • [Benson]: Update when automated handling of multiple spectral windows is available (Q2). *Clarified: John will organize a small group this week (likely necessary candidates are Eric, Michael, Rick) to resolve the near and long term plans for handling the automated export to UVFITS for heterogeneous spectral windows (e.g., auto split by spectral window, etc). Results to be presented next Monday. Bryan - will lead this.*


  • Dynamic scheduler use of API (saturation phase vs. rms phase); use must be consistent everywhere. Frazer noted that we need configuration-based criteria.

Planning review (for Q1/Q2 2010):

Note: Beginnings of group pages from the organization page:

Followup deferred. Begin to look at the detailed needs for the tests being scoped out by the leads. Want to finalize the detailed observing/testing plans by the end of this month and have a view of the longer term commitments needed. Establish an overview of the obstacles to deploying to the community; in particular, understand all of the details of the end-to-end data flow/processing to make decisions on what is working versus what is usable by the community. Something like:

Mode OPT Scheduling Executor/Corr Calibration strategy Archive CASA AIPS Data Distribution Other Overall Status
Example limit of 1 s int ok OPT selected limits not possible no recommendations avail ok can't cook coffee weather table - transient issue no-go

  • Critical Commissioning
    • System Verification (Perley/Rupen): Need plan to conclude complex gain, bandpass stability tests at L-Q, reference pointing, gain curves, tipping scans.
    • Observing Band Checkout (principally calibration strategies, flagging)
    • Observing Mode Support (broad modes based on science goals)
    • Observing support (e2e capability)
      • OSRO Testing (Gustaaf); ECSV-18 - needs to be pushed into JIRA.
      • Flux Calibrator Models (Amy); ECSV-43
  • Advanced Science Support

    • Need:
      • [Butler] Planning: Planetary observing (OPT setup; correlator needs); ECSV-7, ECSV-15
      • [Chandler] Planning: Demo Science
      • [Frail] Planning: ToO scheduling heuristics
      • [Sjouwerman] Planning: OPT tests for primary frequencies; ECSV-6
      • [van Moorsel] Planning: Data distribution plan
      • [Sjouwerman] Planning: Archive functionality tests; ECSV-11
      • [Mioduszewski] Planning: Calibrator survey; ECSV-27
      • [Brisken] Planning: Phased array (VLBI, pulsars); ECSV-22, ECSV-23
      • [Golap] Planning CASA multi-thread/cluster testing.
      • [Walker] Planning VLBI setup
-- JosephMcMullin - 26 Apr 2010
Topic revision: r6 - 2010-04-28, jmcmulli
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback